# Efficient Online Inference of Vision Transformers by Training-Free Tokenization

### v1

@@ C.3 Vocabulary Size
${\mathcal{T}}_{inter}^{100}$ 66.60 $\pm$ 0.28 9.61 $\pm$ 1.51 90.24 $\pm$ 0.06 56.48 $\pm$ 4.32 76.16 $\pm$ 0.70 61.43 $\pm$ 1.32 41.06 39.84
${\mathcal{T}}_{inter}^{1000}$ 78.10 $\pm$ 0.18 22.74 $\pm$ 1.24 90.25 $\pm$ 0.15 47.96 $\pm$ 2.24 94.20 $\pm$ 0.63 87.18 $\pm$ 1.78 70.29 63.48
${\mathcal{T}}_{inter}^{10000}$ 78.98 $\pm$ 0.07 21.96 $\pm$ 0.31 89.61 $\pm$ 0.07 47.46 $\pm$ 2.62 95.27 $\pm$ 0.97 87.18 $\pm$ 2.35 70.52 63.40
(b) CLIP (image classification and group robustness)
@@
Table 6: Token length (including [CLS]) and performance of VWTs with random matching. The pairwise cosine distance is initialized by sampling from a uniform distribution of $[0,2]$. Average and worst-group accuracies degrade noticeably with ${\mathcal{T}}_{100}$ except for CelebA. Performance does not change much for ${\mathcal{T}}_{1000}$ and ${\mathcal{T}}_{10000}$ from $Base$ since barely any compression occurs.
@@
Waterbirds 0.2 97 76.71 $\pm$ 0.24 24.09 $\pm$ 0.59
0.3 79 74.85 $\pm$ 0.15 21.13 $\pm$ 0.36
0.4 66 73.35 $\pm$ 0.12 16.30 $\pm$ 0.24
0.5 58 72.81 $\pm$ 0.10 13.14 $\pm$ 0.48
CelebA 0.2 66 89.44 $\pm$ 0.08 56.48 $\pm$ 1.70
0.3 55 88.85 $\pm$ 0.26 58.70 $\pm$ 4.21
0.4 50 88.31 $\pm$ 0.14 57.04 $\pm$ 3.26
0.5 48 88.21 $\pm$ 0.32 55.74 $\pm$ 1.40
MetaShift 0.2 84 86.16 $\pm$ 0.89 77.46 $\pm$ 2.34
0.3 69 81.16 $\pm$ 1.00 67.77 $\pm$ 1.43
0.4 60 77.73 $\pm$ 1.03 60.65 $\pm$ 2.57
0.5 55 74.75 $\pm$ 1.36 56.08 $\pm$ 3.38
@@
Table 7: Token length (including [CLS]) and performance using ${\mathcal{T}}_{inter}^{100}$ with varying similarity thresholds. Higher thresholds can reduce sequences to 48 tokens on CelebA. Greater compression results in higher degradation in performance, particularly in robustness.
Model Subgroup Waterbirds CelebA MetaShift
Length $\Delta$ Accuracy Length $\Delta$ Accuracy Length $\Delta$ Accuracy
$Base$ 0 197 98.89 197 95.63 197 98.54
1 82.45 96.23 93.19
2 21.86 48.67 87.69
3 54.73 50.00 95.36
${\mathcal{T}}_{inter}^{100}$ 0 144 -0.67 85 -3.25 118 -0.74
1 106 -3.82 88 -0.13 110 -6.18
2 140 19.00 98 32.62 119 -2.92
3 105 6.17 99 6.30 100 -9.02
${\mathcal{T}}_{inter}^{1000}$ 0 160 -0.40 119 -0.46 138 -0.25
1 129 -0.22 117 0.74 138 -2.06
2 160 9.03 124 18.92 143 -1.17
3 129 2.18 124 -4.07 131 -2.13
${\mathcal{T}}_{inter}^{10000}$ 0 180 -0.09 158 0.25 163 -0.37
1 156 1.34 151 0.66 167 0.19
2 181 4.75 158 2.96 169 -1.75
3 158 1.71 156 -6.30 163 -0.81
Table 8: Distribution of token length (including [CLS]) and accuracy (w.r.t. $Base$) by subgroup. Tokenizers of varying vocabulary sizes are shown. Like text tokenizers, VWTs may induce significantly unequal token lengths as seen with ${\mathcal{T}}_{inter}^{100}$ on Waterbirds. Subgroups are also affected unequally as stronger compression does not correlate with greater improvement or degradation in performance.
(a) Waterbirds
(b) CelebA
(c) MetaShift
Figure 6: Probability distribution of the matched visual words. Tokenizers of varying vocabulary sizes are shown. The probability distribution exhibits a large skew irrespective of the dataset as certain visual words are matched more frequently than others. Larger vocabularies display greater sparsity as the many visual words that remain unmatched may be pruned.
Figure 7: Visualization of long-form captions by $Q_{8}$, $ToME$, and ${\mathcal{T}}_{inter}^{10000}$ on COCO.
Dataset Model ${\mathcal{T}}_{intra}^{0.25}$ ${\mathcal{T}}_{intra}^{0.33}$ ${\mathcal{T}}_{intra}^{0.7}$
Waterbirds CLIP 148 132 59
CelebA
MetaShift
OpenImages (Com.)
OpenImages (Rare)
COCO BLIP 433 386 173
NoCaps
(a) Length
Model Waterbirds CelebA MetaShift OpenImages
${\mathcal{T}}_{intra}^{0.25}$ 78.31 $\pm$ 0.09 26.69 $\pm$ 0.24 89.46 $\pm$ 0.16 48.47 $\pm$ 1.56 94.70 $\pm$ 0.43 87.69 $\pm$ 1.54 69.79 63.37
${\mathcal{T}}_{intra}^{0.33}$ 77.72 $\pm$ 0.24 27.78 $\pm$ 0.48 89.31 $\pm$ 0.22 48.83 $\pm$ 1.42 94.62 $\pm$ 0.41 87.18 $\pm$ 2.35 69.11 62.75
${\mathcal{T}}_{intra}^{0.7}$ 73.27 $\pm$ 0.12 18.80 $\pm$ 0.32 89.19 $\pm$ 0.05 45.37 $\pm$ 1.28 82.61 $\pm$ 1.10 63.59 $\pm$ 0.89 46.83 42.68
(b) CLIP (image classification and group robustness)
Model COCO NoCaps
BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
${\mathcal{T}}_{intra}^{0.25}$ 33.92 107.47 103.59 14.65 98.39 14.07 94.47 13.59 98.34 14.06
${\mathcal{T}}_{intra}^{0.33}$ 33.59 106.25 103.19 14.65 97.40 13.95 92.66 13.41 97.27 13.95
${\mathcal{T}}_{intra}^{0.7}$ 29.64 93.20 89.89 13.53 85.04 12.82 80.81 12.31 84.88 12.82
(c) BLIP (image captioning)
Table 9: Token length (including [CLS]) and performance of VWTs with varying dropping ratios.
Figure 8: Visualization of patch dropping by our visual word tokenizer (${\mathcal{T}}_{intra}^{0.5}$ model). Patches with the lowest pixel variance that are dropped are indicated in black. In most cases, the dropped patches correspond to the background object that is typically uninformative.
Figure 9: Visualization of captions by token dropping on COCO. Token dropping of up to 0.5 may not cause significant deviation in the generations from the base model. On certain samples, the accuracy of the descriptions can improve as well.
Model Waterbirds CelebA MetaShift OpenImages
${\mathcal{T}}_{inter}^{100}$ 78.41 $\pm$ 0.09 26.01 $\pm$ 0.41 90.04 $\pm$ 0.22 53.15 $\pm$ 3.21 90.24 $\pm$ 0.40 84.28 $\pm$ 2.38 62.90 58.10
${\mathcal{T}}_{inter}^{10000}$ 79.68 $\pm$ 0.11 22.90 $\pm$ 0.16 90.12 $\pm$ 0.18 46.85 $\pm$ 4.63 94.81 $\pm$ 0.95 86.15 $\pm$ 1.54 69.03 62.50
(a) CLIP (image classification and group robustness)
Model COCO NoCaps
BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
(b) BLIP (image captioning)
Table 10: Token length (including [CLS]) and performance of VWTs with varying vocabulary sizes.
@@
Figure 10: Visualization of captions by VWTs on COCO. Increasing the similarity threshold causes the generated captions to deviate more than reducing the vocabulary size. With random matching, the model begins to completely misunderstand the image.

### v2

@@ 4.4 VWTs and Visual Performance
${\mathcal{T}}_{inter}^{100}$ 66.60 9.61 90.24 56.48 76.16 61.43 41.06 39.84
${\mathcal{T}}_{inter}^{1000}$ 78.10 22.74 90.25 47.96 94.20 87.18 70.29 63.48
${\mathcal{T}}_{inter}^{10000}$ 78.98 21.96 89.61 47.46 95.27 87.18 70.52 63.40
(b) CLIP (image classification and subgroup robustness)
@@
Table 5: Token length (including [CLS]) and performance of the inter-image approach with random merging. The pairwise cosine distance is initialized by sampling from a uniform distribution of $[0,2]$. Average and worst-group accuracies degrade noticeably with ${\mathcal{T}}_{inter}^{100}$ except for CelebA. Performance does not change much for ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ from $Base$ since barely any compression occurs.
4.5 Visualization of the Visual Words
To better understand the VWT, we visualize the patch matching of ${\mathcal{T}}_{intra}^{0.5}$ and ${\mathcal{T}}_{inter}^{100}$ in Figures 3 and 1, respectively. For the latter, we highlight in identical colors patches that are matched with one another. With ${\mathcal{T}}_{intra}^{0.5}$, the patches with the lowest pixel variance typically correspond to uninformative backgrounds. In most cases, the dropping of such patches continues to preserve the foreground object.
With ${\mathcal{T}}_{inter}^{100}$, we find that patches representing the background are more frequently grouped than those of the foreground object. On Waterbirds, the merging of background patches may explain the improved robustness in Table 3 by mitigating the spurious correlation with the background. On CelebA, ${\mathcal{T}}_{inter}^{100}$ tends to avoid matching the eyes or mouths of the individuals. We also observe that patch matching is capable of grouping similar but non-adjacent visual concepts. In certain examples of MetaShift and NoCaps, multiple cats and foods are seemingly matched together, respectively. Our analysis shows that patch matching serves as a rudimentary form of image segmentation by identifying similar visual concepts for compression.
In Figure 4, we visualize the vocabulary of ${\mathcal{T}}_{inter}^{100}$ to analyze the formation of the visual words. We show patches from ImageNet-1K (i.e. pre-processing data) that are matched to each visual word using the Euclidean distance. Since visual words are centroids, patches that are matched to the same visual word belong to the same cluster. We observe that visual words depict basic features such as colors or edges. These features are also formed as an average representation of the patches that belong to each cluster. By representing basic features, visual words serve as effective anchor points to which patches can be matched to.
@@ 4.5 Visualization of the Visual Words
Waterbirds 0.2 97 76.71 24.09
0.3 79 74.85 21.13
0.4 66 73.35 16.30
0.5 58 72.81 13.14
CelebA 0.2 66 89.44 56.48
0.3 55 88.85 58.70
0.4 50 88.31 57.04
0.5 48 88.21 55.74
MetaShift 0.2 84 86.16 77.46
0.3 69 81.16 67.77
0.4 60 77.73 60.65
0.5 55 74.75 56.08
@@
Table 6: Token length (including [CLS]) and performance using ${\mathcal{T}}_{inter}^{100}$ with varying similarity thresholds. Higher thresholds can reduce sequences to 48 tokens on CelebA. Greater compression results in higher degradation in performance, particularly in subgroup robustness.
4.6 Ineffectiveness of the Embedding Space
In Table 4, we analyze the sequence compression using visual words formed in the embedding space. Instead of initializing the vocabulary by clustering patches in the pixel space, we do so with patches after the pre-trained convolution layer of CLIP or BLIP. During inference, we match the patches after the pre-trained convolution layer with these new visual words. Compared to Table 1, we observe a notable reduction in the compression irrespective of pre-processing data and vocabulary size. In Tang et al. [55], embeddings are shown to become progressively more similar to one another due to self-attention. As such, it is unsurprising that matching patches and visual words in the initial embedding space is found to be ineffective.
4.7 Random Merging of Tokens
In Table 5, we study the effects of randomly merging the tokens. We initialize the pairwise cosine distance by sampling from a uniform distribution of $[0,2]$. First, we find the token length (including [CLS]) to differ noticeably from Table 1. For ${\mathcal{T}}_{inter}^{100}$, sequences are further reduced across the datasets. Conversely, ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ display little compression. Second, performance is shown to change from Table 3. We observe a significant degradation in average and worst-group accuracies with ${\mathcal{T}}_{inter}^{100}$ on Waterbirds and MetaShift. For ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$, performance does not shift much from $Base$ as barely any compression occurs. With random merging, the captions are shown to deviate completely from those of $Base$ in Figure 6, thus further demonstrating that visual words are an effective criterion for grouping patches as shown in Figure 1.
4.8 Ablation of the Similarity Threshold
We have shown how the inter-image approach can improve the efficiency of online inference. To better understand their limitations, we ablate the threshold of ${\mathcal{T}}_{inter}^{100}$ by setting it to values of $\{0.2,0.3,0.4,0.5\}$ in Table 6. We seek to determine if exploiting higher thresholds for increased compression is a viable strategy. Naturally, we observe a reduction in performance as increasingly dissimilar patches are merged. For Waterbirds and MetaShift, the WGA degrades more significantly than the average, especially with the former. Interestingly, average accuracy remains relatively unchanged while WGA improves significantly on CelebA irrespective of similarity threshold. We posit that at higher thresholds, the merging of core features represented by the foreground object results in the reduced performance of Waterbirds and MetaShift.
@@ 4.8 Ablation of the Similarity Threshold
Figure 6: Visualization of image captions on COCO by the inter-image approach. The generated captions are shown to deviate more when increasing the similarity threshold than when reducing the vocabulary size. With random matching, the model begins to completely misunderstand the image.
4.9 When to use the intra-image or inter-image approach?
Concerning the choice between the intra-image or inter-image approach, we posit that when global information is required, dropping tokens may be more advantageous by removing unnecessary noise from the visual input. On the other hand, when the task necessitates local information (e.g. long-form captioning), merging tokens may better preserve the visual concepts. For example, the image on the first row of Figure 7 of Appendix 6.2 shows that ${\mathcal{T}}_{intra}^{0.5}$ removes the mountainous background, thus leading to the absence of ”mountain” in the long-form caption of Figure 5. Potentially, combining both approaches may maximize compression while preserving information by merging the mountains and dropping the bushes.
5 Conclusion
In this work, we set out to define a training-free tokenization for ViTs that lowers wattage while balancing costs to runtime and performance. In online scenarios, we have shown empirically that our intra-image and inter-image approaches are stronger than 8-bit quantization and token merging for image classification and captioning. Analysis on large-scale classification further validates the viability of our method while long-form captioning shows its potential for improving descriptiveness and coherence. Qualitatively, we observe how the criterion of the intra-image approach typically corresponds to the background while that of the inter-image approach groups analogous visual concepts based on visual words that represent basic features. As a future work, additional research could explore combining both the intra-image and inter-image approaches, potentially improving further the performance and efficiency of the VWT.
Acknowledgements